OpenStudy (anonymous):

the sum of a number and 2 is 6 less than twice that number.

OpenStudy (anonymous):

the sum of a number and 2 is 6 less than twice that number.

OpenStudy (pixiedust1):

The sum of a number and 2 is 6 less than twice that number. Let's break this down. The sum of a number and 2 Sum is addition n+2 is means = n+2= 6 less than twice that number less than is - 6 - twice that number is 2n (twice of n) 6-2n Put that all together and n+2=6-2n

OpenStudy (pixiedust1):

The sum of a number and 2 is 6 less than twice that number. n+2=6-2n Its okay if you're having trouble with this concept, its difficult. I struggled with it for months, but now i fully understand.